# Multi-target quantum compilation algorithm

### v1

@@ 4.3 Detailed experimental setting in the Results
and get the best circuits for each setup. We use these best circuits to calculate the risk on the testing set. All results in Fig. 2 are for the testing set.
Thermal state preparation
Dense-purification method. We examine thermal state preparation using the dense-purification method across 2 to 10 qubits. Parameters include $n_{\rm iter}=100$, $d=2N$, $n_{\mathcal{V}}=8$, $n_{\rm gene}=16$. Note that $n_{\rm gene}$ does not significantly affect the runtime, as the program can stop anytime if the threshold is met. The threshold and optimizer remain consistent with the previous settings.
Conventional purification method. For the conventional purification method, we examine for $N=2$ thermal state, which requires a 4-qubit circuit. Initially, we generate 100 target TFD states corresponding to equally-spaced $\beta\in[0,10]$. We then run GA-VQA to find the best circuit structure for all those TFD states based on the weighted-sum cost function as we will explain below. Finally, we continue to run optimization using Adam optimizer for 100 iterations to find optimal parameters for each quantum state.
@@
$\mathcal{F}_{\rm ws}=\frac{w_{1}*\mathcal{F}_{\beta\in[0,4)}+w_{2}*\mathcal{F}_{\beta\in[4,7)}+w_{3}*\mathcal{F}_{\beta\in[7,10]}}{w_{1}+w_{2}+w_{3}},$ (15)
@@
Time-dependent dynamic simulation
GA-VQA method. For the GA-VQA method, our simulation follows these steps: (1) Create the initial state $|\psi_{0}\rangle=|11...00\rangle$ by applying X gates to $N/2$ qubits. (2) Create a set of 100 target unitaries $U(t_{i})$ using Eq. (11) for $t_{i}\in[0,10],\ i=\{1,2,\cdots,100\}$. (3) We then run GA-VQA to find the best circuit structure for all those target unitaries. Once $V^{*}(\bm{\Theta}^{*})$ is determined, we measure the magnetization as follows. (i) Prepare $|\psi_{0}\rangle$ in a quantum circuit. (ii) Apply $V^{*}(\bm{\Theta}^{*})$ to get the final state $|\check{\psi}(t)\rangle=V^{*}(\bm{\Theta}^{*})|\psi_{0}\rangle$. (iii) Measure qubit $j$ and get the probability
$p_{j}(m)=\langle\check{\psi}(t)|\Pi_{j}|\check{\psi}(t)\rangle,\text{ for }m\in\{0,1\},$ (16)
$M_{j}(t)=p_{j}(0)-p_{j}(1).$ (17)
Trotterization method. To implement the trotterization, we break down the evolution $U(t_{i})$ as follows
[…]
$=\exp\Bigg{\{}-i\frac{\delta t}{K}\lim_{K\to\infty}\Big{[}H\big{(}t_{i-1}\big{)}+H\big{(}t_{i-1}+\frac{\delta t}{K}\big{)}+\cdots+H\big{(}t_{i-1}+\frac{(K-1)\delta t}{K}\big{)}\Big{]}\Bigg{\}},$ (18)
where $\delta t=t_{i}-t_{i-1}=0.1$ is the interval, divided into $K$ equal sub-intervals. Here, we set $K=5$, making $\delta t/K=0.02$ sufficiently small. Each $H(s)$ in the above expression consists of single and two-qubit gates, and we assume $H(s)=\sum_{j=1}^{L}H_{j}$. Using trotterization, we can break $H(s)$ own into a sequence of quantum gates and implement it in quantum circuits. The trotterization expansions of the first order (Trotter1) and second order (Trotter2) are given by [50, 44]
[…]
where $r$ is the Trotter number. In this work, we set $r=100$.
AVQDS method. The AVQDS method, introduced in Ref. [7], works as follows. Starting with the initial quantum state $|\psi_{0}\rangle$, the state evolves over time according to $|\psi(t)\rangle$ as
$\dfrac{d|\psi(t)\rangle}{dt}=-iH(t)|\psi(t)\rangle.$ (21)
To solve this, we use a variational quantum ansatz $|\psi(\bm{\theta}(t)\rangle$, where $\bm{\theta}(t)$ represents a set of time-dependent parameters. For example, at time $t=0$, it gives $|\psi(\bm{\theta}(t=0))\rangle=|\psi_{0}\rangle$. These parameters are trained to minimize the squared McLachlan distance [51] $\mathcal{L}^{2}$
$\mathcal{L}^{2}=\Big{\|}\sum_{\mu}\dfrac{d|\psi(\bm{\theta})\rangle}{d\theta_{\mu}}\dfrac{d\theta_{\mu}}{dt}+iH(t)|\psi(\bm{\theta})\rangle\Big{\|}^{2}_{F},$ (22)
In our numerical simulation, we first construct the Hamiltonian $H(t)$ as described in Eq. (10). We then employ the pseudo-Trotter ansatz [7] $|\psi(\bm{\theta})\rangle=\prod_{\mu=1}^{N_{\bm{\theta}}}e^{-i\theta_{\mu}A_{\mu}}|\psi_{0}\rangle$, where $N_{\bm{\theta}}$ is the number of trainable time-dependent parameters, and $A_{\mu}$ are Hermitian operators from a set of Pauli operators. During the simulation for any time $0<t_{i}<T$, the ansatz is adaptively updated until the squared McLachlan distance meets the threshold $\mathcal{L}^{2}_{\rm cut}$. We set the threshold $\mathcal{L}^{2}_{\rm cut}=10^{-3}$ for Fig. 4a, and $\mathcal{L}^{2}_{\rm cut}=10^{-1}$ for Fig. 4b,c. In Fig. 4b,c, the AVQDS method is not trainable at $\mathcal{L}^{2}_{\rm cut}=10^{-3}$, so we temporarily reduce its accuracy by training at $\mathcal{L}^{2}_{\rm cut}=10^{-1}$. This is why the accuracy of this method is lower compared to the exact resutl and GA-VQA method.
Variational quantum eigensolver
For the variational quantum eigensolver [46, 52], an upper bound of ground state energy $E_{o}$ of a given Hamiltonian is bounded by
$E_{o}\leq\frac{\langle\psi(\theta)|H|\psi(\theta)\rangle}{\langle\psi(\theta)|\psi(\theta)\rangle}\equiv\langle H\rangle,$ (23)
which is found by optimizing the parameters of quantum state $|\psi\rangle\equiv V_{l}(\theta)|\textbf{0}\rangle$. To simulate the GA-VQA scheme’s application in VQE with a Hydrogen molecule, we run GA-VQA with the setup that can be seen in the VQE field at Tab. 3. In this case, the target of the GA-VQA is to find ansatzes with low depth but still gain the needed accuracy of ground state molecular energy. The Genetic Algorithm is leveraged to find the friendly NISQ ansatz while the optimizing part is left to VQE. This method iteratively converges to higher fidelity (the higher the fidelity, the more accurate the result) until the max generation or termination condition is satisfied.
@@
Data availability
Code availability
The code is available at https://github.com/vutuanhai237/GA-QAS.
Acknowledgment
V.T.H. thanks Dr. Luong Ngoc Hoang for advice on genetic algorithms. This research is funded by Vietnam National University Ho Chi Minh City (VNU-HCM) under grant number C2024-28-04, JSPS KAKENHI Grant Number 23K13025, Unitary Fund, and Tohoku University FRIS URO.
Author contributions statement
Competing interests
The author declares no competing interests.

### v2

@@ B.1 Numerical benchmarking
and get the best circuits for each setup. We use these best circuits to calculate the risk on the testing set. All results in Figure  2 are for the testing set.
B.2 Thermal state preparation
B.2.1 Conventional purification method.
For the conventional purification method, we examine for $N=2$ thermal state, which requires a 4-qubit circuit. Initially, we generate 100 target TFD states corresponding to equally-spaced $\beta\in[0,10]$. We then run GA-VQA to find the best circuit structure for all those TFD states based on the weighted-sum cost function as we will explain below. Finally, we continue to run optimization using Adam optimizer for 100 iterations to find optimal parameters for each quantum state.
@@ B.2 Thermal state preparation
$\mathcal{F}_{\rm ws}=\frac{w_{1}*\mathcal{F}_{\beta\in[0,4)}+w_{2}*\mathcal{F}_{\beta\in[4,7)}+w_{3}*\mathcal{F}_{\beta\in[7,10]}}{w_{1}+w_{2}+w_{3}},$ (B.2)
@@
B.2.2 Dense-purification method.
We examine thermal state preparation using the dense-purification method across 2 to 8 qubits. Parameters include $n_{\rm iter}=100$, $d=2N$, $n_{\mathcal{V}}=8$, $n_{\rm gene}=16$. Note that $n_{\rm gene}$ does not significantly affect the runtime, as the program can stop anytime if the threshold is met. The threshold and optimizer remain consistent with the previous settings. Note that in this case, the cost function is fidelity $\mathcal{F}$ in Eq. (7).
@@
B.3 Time-dependent dynamic simulation
B.3.1 GA-VQA method.
For the GA-VQA method, our simulation follows these steps: (1) Create the initial state $|\psi_{0}\rangle=|11...00\rangle$ by applying X gates to $N/2$ qubits. (2) Create a set of 100 target unitaries $U(t_{i})$ using Eq. (9) for $t_{i}\in[0,10],\ i=\{1,2,\cdots,100\}$. (3) We then run GA-VQA to find the best circuit structure for all those target unitaries. Once $V^{*}(\bm{\Theta}^{*})$ is determined, we measure the magnetization as follows. (i) Prepare $|\psi_{0}\rangle$ in a quantum circuit. (ii) Apply $V^{*}(\bm{\Theta}^{*})$ to get the final state $|\check{\psi}(t)\rangle=V^{*}(\bm{\Theta}^{*})|\psi_{0}\rangle$. (iii) Measure qubit $j$ and get the probability
$p_{j}(m)=\langle\check{\psi}(t)|\Pi_{j}|\check{\psi}(t)\rangle,\text{ for }m\in\{0,1\},$ (B.3)
$M_{j}(t)=p_{j}(0)-p_{j}(1).$ (B.4)
B.3.2 Trotterization method.
To implement the Trotterization, we break down the evolution $U(t_{i})$ as follows
[…]
$=\exp\Bigg{\{}-i\frac{\delta t}{K}\lim_{K\to\infty}\Big{[}H\big{(}t_{i-1}\big{)}+H\big{(}t_{i-1}+\frac{\delta t}{K}\big{)}+\cdots+H\big{(}t_{i-1}+\frac{(K-1)\delta t}{K}\big{)}\Big{]}\Bigg{\}},$ (B.5)
where $\delta t=t_{i}-t_{i-1}=0.1$ is the interval, divided into $K$ equal sub-intervals. Here, we set $K=5$, making $\delta t/K=0.02$ sufficiently small. Each $H(s)$ in the above expression consists of single and two-qubit gates, and we assume $H(s)=\sum_{j=1}^{L}H_{j}$. Using Trotterization, we can break $H(s)$ own into a sequence of quantum gates and implement it in quantum circuits. The Trotterization expansions of the first order (Trotter1) and second order (Trotter2) are given by PhysRevLett.128.210501 ; ikeda2023trotter24
[…]
where $r$ is the Trotter number. In this work, we set $r=100$.
B.3.3 AVQDS method.
The AVQDS method, introduced in Ref. PRXQuantum.2.030307 , works as follows. Starting with the initial quantum state $|\psi_{0}\rangle$, the state evolves over time according to $|\psi(t)\rangle$ as
$\dfrac{d|\psi(t)\rangle}{dt}=-iH(t)|\psi(t)\rangle.$ (B.8)
To solve this, we use a variational quantum ansatz $|\psi(\bm{\theta}(t)\rangle$, where $\bm{\theta}(t)$ represents a set of time-dependent parameters. For example, at time $t=0$, it gives $|\psi(\bm{\theta}(t=0))\rangle=|\psi_{0}\rangle$. These parameters are trained to minimize the squared McLachlan distance doi:10.1080/00268976400100041 $\mathcal{L}^{2}$
$\mathcal{L}^{2}=\Big{\|}\sum_{\mu}\dfrac{d|\psi(\bm{\theta})\rangle}{d\theta_{\mu}}\dfrac{d\theta_{\mu}}{dt}+iH(t)|\psi(\bm{\theta})\rangle\Big{\|}^{2}_{F},$ (B.9)
In our numerical simulation, we first construct the Hamiltonian $H(t)$ as described in Eq. (8). We then employ the pseudo-Trotter ansatz PRXQuantum.2.030307 $|\psi(\bm{\theta})\rangle=\prod_{\mu=1}^{N_{\bm{\theta}}}e^{-i\theta_{\mu}A_{\mu}}|\psi_{0}\rangle$, where $N_{\bm{\theta}}$ is the number of trainable time-dependent parameters, and $A_{\mu}$ are Hermitian operators from a set of Pauli operators. During the simulation for any time $0<t_{i}<T$, the ansatz is adaptively updated until the squared McLachlan distance meets the threshold $\mathcal{L}^{2}_{\rm cut}$. We set the threshold $\mathcal{L}^{2}_{\rm cut}=10^{-3}$ for Figure  4(a), and $\mathcal{L}^{2}_{\rm cut}=10^{-1}$ for Figure  4(b,c). In Figure  4(b,c), the AVQDS method is not trainable at $\mathcal{L}^{2}_{\rm cut}=10^{-3}$, so we temporarily reduce its accuracy by training at $\mathcal{L}^{2}_{\rm cut}=10^{-1}$. This is why the accuracy of this method is lower compared to the exact result and GA-VQA method.
B.4 Variational quantum eigensolver
For the variational quantum eigensolver Peruzzo2014 ; TILLY20221 , an upper bound of ground state energy $E_{o}$ of a given Hamiltonian is bounded by
$E_{o}\leq\frac{\langle\psi(\theta)|H|\psi(\theta)\rangle}{\langle\psi(\theta)|\psi(\theta)\rangle}\equiv\langle H\rangle,$ (B.10)
which is found by optimizing the parameters of quantum state $|\psi\rangle\equiv V_{l}(\theta)|\textbf{0}\rangle$. To simulate the GA-VQA scheme’s application in VQE with a Hydrogen molecule, we run GA-VQA with the setup that can be seen in the VQE field at Tab. 3. In this case, the target of the GA-VQA is to find ansatzes with low depth but still gain the needed accuracy of ground state molecular energy. The GA is leveraged to find the friendly NISQ ansatz while the optimizing part is left to VQE. This method iteratively converges to higher fidelity (the higher the fidelity, the more accurate the result) until the max generation or termination condition is satisfied.